# ReVision: High-Quality, Low-Cost Video Generation with Explicit 3D Physics Modeling for Complex Motion and Interaction

### v1

@@ 4.2 Proposed Method: ReVision
Overview. As shown in Fig. 2, ReVision consists of three stages. In stage one (S1), we generate a coarse video based on the provided conditions. In stage two (S2), we extract both 2D and 3D features from the coarse video and refine the 3D motion sequences through the proposed PPPM. In stage three (S3), we use the refined 3D motion sequences as strong conditioning, guiding the video generation model to regenerate the video, resulting in significantly higher-quality output even for complex motions and interactions.
@@
S2: Object-Centric 3D Optimization. Following coarse video generation, we utilize 3D parametric mesh models [32, 49, 80] for humans and animals, along with 2D bounding boxes [57], segmentation masks [71, 40], and depth estimation [69] for general objects, to parameterize targets in 3D space. These parameters are extracted using a series of off-the-shelf models (see Sec. 3). However, due to the poor motion quality and inconsistencies in the coarse video generated in S1, the 3D parameters extracted also suffer from instability and inconsistencies. To address this, we propose a Parameterized Physical Prior Model (PPPM) to optimize the 3D motion sequence, based on the text information and the motion strength.
@@
PPPM first extracts text embeddings from the text description using a pre-trained CLIP encoder [42]. Motion strength is computed from the differences in parametric 3D model parameters between adjacent frames, providing a measure of motion speed. Since the 3D motion sequences are already parameterized as 2D vectors, PPPM employs a series of transformer blocks to iteratively refine the motion sequence based on these conditioning inputs. Within each block, motion features undergo self-attention, followed by cross-attention with the conditioning inputs (text embeddings and motion strength) and a feedforward network to generate the final output. Finally, the optimized 3D parameterized motion sequences are converted into 3D mesh sequences and projected into 2D as part segmentation masks and confidence maps, providing more accurate motion guidance. Architectural details can be found in Sec. 9 of Supp.
@@
To effectively train PPPM, we introduce small perturbations to ground-truth motion sequences during training. Three types of perturbations are randomly applied: (1) adding small random noise to the motion sequence, (2) shuffling the internal order of the sequence, and (3) dropping a small segment while repeating the remaining segments to maintain the original length. Through this process, PPPM learns to denoise perturbations, improving its ability to recover smooth and robust motion sequences.
@@
5 Dataset
Figure 4: Qualitative results. ReVision enables the generation of high-quality videos with complex motions and interactions, such as running with dogs, picking up a ball, and hitting a tennis ball. It allows precise motion control over small body parts, such as arms and hands. Zoom in for better detail. Please check the videos on the project page.
Both the motion-conditioned video generation model and the parameterized physical prior model need to be fine-tuned (trained) on a small yet high-quality video dataset with object-centric annotations. Existing large-scale video datasets [2, 8] mainly provide text-image pairs without detailed object-centric annotations. To address this limitation, we use a suite of off-the-shelf models across various tasks to generate detailed 2D and 3D object-centric annotations. We annotate a total of 20K videos from the Panda-70M [8] dataset. For each video, we provide frame-wise 2D bounding boxes, semantic masks, depth estimation maps, and 3D parametric mesh reconstructions for detected humans and animals. The details are outlined below.
High-Quality Motion Videos Filtering. To start with, we use LLMs [68] and an open-vocabulary segmentation model [71] to curate high-quality motion videos. Specifically, the language model filters videos with evident motion based on their captions. Then, for each selected video, we equally sample 10 frames and apply the segmentation model to identify humans and animals. We evaluate each frame based on the predicted mask size and mask count. Then we retain videos where humans or animals occupy a significant portion of the frame and where the count of humans does not exceed five in each frame.
Object Detection and Depth Estimation. Based on the captions of the selected videos, we identify the objects mentioned and detect their bounding boxes [57] and instance masks [71]. Additionally, we apply Depth Anything V2 [69] to generate the depth maps of each frame.
Human Videos Annotation. For videos containing humans, we focus on extracting 2D instance segmentation masks, 2D part masks, 2D face keypoints, 3D body pose and shape, and 3D hand pose. We begin by using YOLO-V8 [57] to segment all humans in each frame, providing accurate human masks. Next, we apply a state-of-the-art face keypoint detector, RTMPose [20], to predict facial keypoints for each detected human. Simultaneously, we use 4D-Human [12] and HaMeR [38] to estimate the 3D body and hand meshes. The resulting SMPL (body mesh) [32] and MANO (hand mesh) [47] parameters are then fit into a unified SMPL-X [37] representation, which contains both human body and hand meshes. We then project the 3D SMPL-X human mesh onto 2D to obtain part masks, as each vertex in the SMPL-X mesh is labeled by body part. Finally, we project the face keypoints and 3D human mesh onto the instance mask, allowing us to compute the overlap between the projected keypoints, projected human mask, and detected 2D human mask. This overlap is quantified using an IoU score, which is used to filter out annotations with high errors. As a result, for each video, we obtain annotations including human instance masks, 2D facial keypoints, 3D SMPL-X meshes for the body and hands, and 2D part-level segmentation masks.
Animal Videos Annotation. We start by using Grounded SAM 2 [23, 43, 44] to segment animal masks in each frame. Next, we apply a state-of-the-art camera estimation algorithm, VGGSfM [59], to optimize the camera’s intrinsic and extrinsic parameters across the video. To ensure a reliable camera estimate, we set thresholds on mean projection errors and mean track lengths, filtering out videos that do not meet these criteria. We then use AnimalAvatar [50] initialized with Animal3D [66] to fit SMAL parameters. Each video is divided into segments of 10 consecutive frames, and AnimalAvatar is applied to each segment independently. This strategy helps mitigate the impact of outliers in camera predictions on the overall optimization quality. To ensure the accuracy of SMAL fitting, we impose thresholds on IoU and PSNR [15], filtering out video segments that do not meet our accuracy standards. Once accurate SMAL fittings are obtained, we follow a similar pipeline to extract the desired annotations as used in human cases.
@@ 6 Experimental Results
In this section, we first compare our method with Stable Video Diffusion (SVD) [6] and HunyuanVideo [24] in Sec. 6.1, demonstrating how our approach enhances SVD to enable more controllable and complex motion generation. Next, in Sec. 6.2, we compare our model with Human Image Animation methods, demonstrating that it achieves comparable performance even when given only the final pose instead of the full motion sequence. Finally, in Sec. 6.3, we conduct ablation studies to validate the effectiveness of our key components. Additional experimental results are provided in the Supp.
@@ 6.1 Image-to-Video Generation
Experimental Setup. We use SVD [6] as the base video generation model and modify it by introducing two additional channels for conditional generation. We fine-tune SVD on our annotated dataset for 300K iterations with a batch size of 64 and a constant learning rate of $2\times 10^{-5}$. During training, we randomly sample 16-frame video clips with a stride of 4 at a resolution of $1024\times 576$. To enable various control, we incorporate different conditioning strategies: $40\%$ of video clips contain accurate part masks for each frame, $30\%$ contain a polygon mask for random parts in the final frame, and the remaining clips have no additional conditioning.
Motion ConsistencyAmount of MotionMotion Reality0020202020404040406060606080808080100100100100User preferenceReVision-SVD (Ours)   SVD
@@
Motion ConsistencyAmount of MotionMotion Reality0020202020404040406060606080808080User preferenceReVision-SVD (Ours)   HunyuanVideo
@@
Figure 5: User preference comparison. Our model enhances the motion generation capability of the pre-trained SVD. It even surpasses HunyuanVideo, a SOTA model with 13B parameters. These results highlight the effectiveness of our model in generating complex motions and interactions.
@@
User Study. To better compare our model with the baseline SVD and HunyuanVideo, we conduct user studies on Amazon MTurk to assess user preferences. Specifically, we generate $500$ text descriptions of humans and animals engaged in daily activities using GPT-4o [18]. For the comparison with SVD, we use GPT-4o to generate five $16:9$ images for each prompt, which are resized to $1024\times 576$ as input. For the comparison with HunyuanVideo, we first use their released model to generate five videos at a resolution of $1280\times 720$ for each prompt, then extract the first frame of each video as the input image for our model to generate the corresponding video. For each image, we generate one video per model using the same random seed ($42$), resulting in a total of $5,000$ video pairs. Each video pair is evaluated by three different users, leading to a total of $15,000$ comparisons. Users are shown two videos side by side, generated by different models, with the order randomized. They are instructed to assess the videos based on Motion Consistency, Amount of Motion, and Motion Realism. The results are reported in Fig. 5. Our model significantly enhances the motion generation capabilities of SVD, producing videos with superior motion quantity, consistency, and realism. Furthermore, it even surpasses HunyuanVideo, a state-of-the-art video generation model with 13B parameters, in terms of motion quality. These results highlight the effectiveness of our model in generating complex motions and interactions.
@@
Qualitative Results. We provide samples of generated videos in Fig. 4. As shown in the figure, our approach, ReVision, produces realistic movements that closely follow user instructions. It also generates high-quality videos that involves complex motions and interactions, such as running with dogs, picking up a ball, and hitting a tennis ball. More visualizations are available in the supplementary videos.
Inference Speed.
@@
Time (s) 36 36 + 5 + 36 411
Table 1: Inference speed. We compare the average time required to generate a 32-frame video for each method.
We compare the inference speed of our model against two baselines in Tab. 1. Due to the limitation of SVD, we use all three models to generate a video clip with 32 frames. Despite the three-step pipeline, the additional 3D detection and motion refinement modules add only 5 seconds to the inference time on a single A100, making it significantly faster than SVD, which requires 36 seconds. More importantly, integrating these modules into SVD significantly improves its ability to generate intricate motions, making it comparable to or superior to state-of-the-art models like HunyuanVideo, which requires an average of 411 seconds.

### v2

@@ 4.2 Proposed Method: ReVision
Overview. As shown in Fig. 2, ReVision consists of three stages. In stage one (S1), we generate a coarse video based on the provided conditions. In stage two (S2), we extract both 2D and 3D features from the coarse video and refine the 3D motion sequences through the proposed PMP. In stage three (S3), we use the refined 3D motion sequences as strong conditioning, guiding the video generation model to regenerate the video, resulting in significantly higher-quality output even for complex motions and interactions.
@@
Although we utilize only 2D and 3D motion features from the coarse video generated in Stage 1, this phase remains foundational, as it is critical for capturing rich motion patterns, intricate object interactions, and authentic camera movements in complex real-world settings. Video generation is inherently a multifaceted task. Beyond producing realistic object appearances, it also requires an understanding of dynamic motion, scene context, coherent camera trajectories, and the diverse interplay of these elements. While training a motion generation model directly for this task is theoretically feasible, it proves challenging in practice. Current state-of-the-art models (Guo et al., 2024; Zhang et al., 2024b; 2023b) are constrained by the lack of large datasets, limiting their capacity to model only simplistic motions, such as human-like activities like running or dancing. Consequently, they struggle to generalize to more complex motions, diverse object interactions, and fail to generate motions that align with realistic camera dynamics and scene context.
Stage 1 mitigates these limitations by directly generating videos and extracting detailed motion patterns, camera trajectories, scene transitions, and meaningful interactions. By leveraging extensive motion priors learned from billions of videos, it constructs a comprehensive sketch of motion and scene structure. This sketch serves as a critical foundation, providing the diversity and realism necessary to produce lifelike, engaging videos. The subsequent stages build on this: Stage 2 refines the motion further, while Stage 3 focuses on generating the video’s visual appearance based on the refined motion feature.
Notice that, because only rough motion features without detailed visual information are needed from the coarse video, we can significantly reduce computational overhead. For instance, our experiments show that the compute time for Stage 1 can be reduced from 36 seconds to 8 seconds by generating the coarse video at a lower resolution (1/4 of the original), with fewer frames (1/2 of the original), and fewer denoising steps (32 vs. 50), while still preserving comparable final video quality. This optimization makes the overall generation process more efficient and cost-effective (See Tab. 3).
S2: Object-Centric 3D Optimization. After generating coarse videos, we parameterize the 3D information in the scenes for further optimization. For humans and animals, we employ well-established 3D parametric mesh models (Loper et al., 2015; Rueegg et al., 2023; Zuffi et al., 2024). For general objects, where no unified 3D representation or well-established modeling approach exists, we construct a parameterized representation by combining 2D bounding boxes (Varghese & Sambath, 2024), segmentation masks (Yu et al., 2023; Peng et al., 2020), and estimated depth (Yang et al., 2024b). Specifically, given the detected bounding box and segmentation mask, we extract a contour from the mask and approximate it with 16 vertices. We then combine these with 4 bounding box corners and the box center, yielding a total of 21 key 2D points. These points are lifted into 3D space using the estimated depth, resulting in a compact point-based representation for each object, denoted as $p_{o}\in\mathbb{R}^{21\times 3}$.
@@
However, due to the poor motion quality and inconsistencies in the coarse video generated in S1, the 3D parameters extracted also suffer from instability and inconsistencies. To address this, we propose a Parameterized Motion Prior model (PMP) to optimize the 3D motion sequence, based on the text information and the motion strength.
@@
PMP first extracts text embeddings from the text description using a pre-trained CLIP encoder (Radford et al., 2021). Motion strength is computed from the differences in parametric 3D model parameters between adjacent frames, providing a measure of motion speed. Since the 3D motion sequences are already parameterized as 3D vectors, PMP employs a series of transformer blocks to iteratively refine the motion sequence based on these conditioning inputs. Within each block, motion features undergo self-attention, followed by cross-attention with the conditioning inputs (text embeddings and motion strength) and a feedforward network to generate the final output. Finally, the optimized 3D parameterized motion sequences are converted into 3D mesh sequences and projected into 2D as part segmentation masks and confidence maps, providing more accurate motion guidance. Architectural details are provided in Sec A of the Appendix, and the effectiveness of PMP is demonstrated in Sec. 5.3.
@@
To train PMP, we introduce small perturbations to the ground-truth motion sequences from the annotated Panda-70M subset. Three types of perturbations are randomly applied: (1) adding small random noise to the motion sequence, (2) shuffling the internal order of the sequence, and (3) dropping a small segment while repeating the remaining segments to maintain the original length. Through this process, PMP learns to denoise perturbations, improving its ability to recover smooth and robust motion sequences.
@@ 5 Experimental Results
Figure 4: Qualitative comparisons. ReVision generates high-quality videos with complex motions and interactions of humans, animals, and general objects. Zoom in for better details. Please find the side-by-side video comparisons in the supplementary video. Reference frames are in the first column.
We first compare our method with SVD (Blattmann et al., 2023) and HunyuanVideo (Kong et al., 2024) in Sec. 5.1, highlighting how it enhances SVD to support more controllable and complex motion generation while maintaining efficiency, effectively handling occlusions, and enabling long video generation. Next, in Sec. 5.2, we compare our model with Human Image Animation methods, demonstrating its ability to generate complex motions. We then evaluate the effectiveness of the proposed Parameterized Motion Prior in Sec. 5.3. Due to space limitations, additional details and results, including ablations on parametric 3D mesh, text prompt, and motion strength, are provided in Sec. C of the Appendix.
@@ 5.1 Image-to-Video Generation
Dataset. Both the motion-conditioned video generation model and the Parameterized Motion Prior model (PMP) need to be fine-tuned (trained) on a small yet high-quality video dataset with object-centric annotations. Existing large-scale video datasets (Bain et al., 2021; Chen et al., 2024) mainly provide text-image pairs without detailed object-centric annotations. To address this limitation, we use a suite of off-the-shelf models across various tasks to generate 2D and 3D object-centric annotations. We annotate a total of 20K videos from the Panda-70M (Chen et al., 2024) dataset. For each video, we provide frame-wise 2D bounding boxes, semantic masks, depth estimation maps, and 3D parametric mesh reconstructions for detected humans and animals. The details are outlined in Sec. B in the Appendix.
@@
Experimental Setup. For most experiments, we use SVD (Blattmann et al., 2023) as the base video generation model and modify it by introducing two additional channels for conditional generation. We fine-tune SVD on our annotated dataset for 300K iterations with a batch size of 64 and a constant learning rate of $2\times 10^{-5}$. During training, we randomly sample 16-frame video clips with a stride of 4 at a resolution of $1024\times 576$. To enable various control, we incorporate different conditioning strategies: $40\%$ of video clips contain accurate part masks for each frame, $30\%$ contain a polygon mask for random parts in the final frame, and the remaining clips have no additional conditioning.
Motion ConsistencyAmount of MotionMotion Reality02020404060608080100100User preferenceReVision-SVD (Ours)  SVD
@@
Motion ConsistencyAmount of MotionMotion Reality02020404060608080User preferenceReVision-SVD (Ours)  HunyuanVideo
@@
Figure 5: User preference comparisons. Our model enhances the motion generation capability of the pre-trained SVD. It even surpasses HunyuanVideo, a SOTA model with 13B parameters. These results highlight the effectiveness of our model in generating complex motions and interactions.
Table 1: Quantitative comparisons on VBench++. We achieve a significantly higher Dynamic Degree while maintaining similar performance across all metrics of consistency, smoothness, and quality.
Model I2V I2V Subject Background Motion Dynamic Imaging
Type Subject Background Consistency Consistency Smoothness Degree Quality
Step-Video-TI2V (Huang et al., 2025) TI2V 98.63% 98.63% 96.02% 97.06% 99.24% 48.78% 70.44%
DynamiCrafter-1024 (Xing et al., 2024) TI2V 98.17% 98.60% 95.69% 97.38% 97.38% 47.40% 69.34%
Gen-4-I2V (Runway, 2026) I2V 97.84% 97.46% 93.23% 96.79% 98.99% 55.20% 70.41%
Magi-1 (Teng et al., 2025) I2V 98.39% 99.00% 93.96% 96.74% 98.68% 68.21% 69.71%
HunyuanVideo-I2V (Kong et al., 2024) I2V 98.53% 97.37% 95.26% 96.70% 99.23% 22.20% 70.1%
Wan2.1-I2V-14B-720P (Wan et al., 2025) I2V 96.95% 96.44% 94.86% 97.07% 97.90% 51.38% 70.44%
ReVision-Wan2.1 (Ours) I2V 98.10% 97.10% 97.06% 97.89% 97.74% 73.67% 72.86%
SVD-XT-1.1 (Blattmann et al., 2023) I2V 97.51% 97.62% 95.42% 96.77% 98.12% 43.17% 70.23%
ReVision-SVD (Ours) I2V 97.94% 98.06% 96.13% 97.89% 98.88% 83.15% 71.48%
Benchmark on General Video Generation. To better evaluate our model on general video generation, we used VBench++ (Huang et al., 2024), which provides a comprehensive, detailed, multi-dimensional assessment of general video generation quality. As our model is designed for image-to-video generation (I2V), we primarily compared it against SVD-XT-1.1 (Blattmann et al., 2023). Moreover, since our model is backbone-agnostic, we integrate our method with a stronger and more recent video diffusion backbone, Wan2.1-I2V-14B-720P (Wan et al., 2025). We train ReVision-Wan2.1 on the same dataset as ReVision-SVD using the same settings. Results are provided in Tab. 1. Our models consistently outperforms the base models on nearly all metrics, particularly in dynamic degree (83.15% vs. 43.17%, 73.67% vs. 51.38%) and various consistency and smoothness measures.
In addition, we compared our method with recent state-of-the-art models for text+image-to-video (TI2V) generation in Tab. 1. We observe that, although these TI2V models use text descriptions to specify motion, the intensity and quality of the resulting movements are not fully controllable. In contrast, our method achieves substantially better performance in generating complex motions, while maintaining high consistency, temporal smoothness, and overall visual quality.
Table 2: Comparisons with motion-conditioned video generation methods on DAVIS. ReVision achieves more accurate motion transfer while preserving object appearance and scene details, resulting in enhanced temporal coherence and reduced visual artifacts.
CoTracker Optical Flow Pixel Subject Background Motion
mIoU $\uparrow$ Error $\downarrow$ MSE $\downarrow$ Consistency $\uparrow$ Consistency $\uparrow$ Smooth $\uparrow$
MotionClone (Ling et al., 2024) 0.72 0.42 0.068 0.75 0.85 0.92
ImageConductor (Li et al., 2025) 0.66 0.64 0.072 0.77 0.88 0.93
Go-with-the-Flow (Burgert et al., 2025) 0.74 0.36 0.053 0.88 0.92 0.98
ReVision-SVD (Ours) 0.80 0.33 0.046 0.96 0.97 0.99
Comparisons with Motion-conditioned Video Generation. Following Go-with-the-Flow (Burgert et al., 2025), we also consider the motion-transfer image-to-video task on DAVIS (Pont-Tuset et al., 2017) and report the results in Tab. 2. We observe that our model consistently outperforms all baselines across all metrics, clearly demonstrating the effectiveness of the proposed method. In particular, our approach yields more accurate motion transfer while better preserving object appearance and scene details, leading to improved temporal coherence and fewer visual artifacts.
In addition, comparing text-conditioned and motion-conditioned video generation shows that explicit motion conditioning plays a dominant role in controlling dynamics, while text conditioning provides only coarse guidance. Although text descriptions effectively convey high-level semantic intent, they struggle to specify fine-grained temporal details and precise motion patterns. As a result, motion conditioning enables more direct and reliable control over dynamic behaviors that are difficult to express through natural language.
@@
User Study. To better compare our model with the baseline SVD and HunyuanVideo, we conduct user studies to assess user preferences. Specifically, we generate $500$ text descriptions of humans and animals engaged in daily activities using GPT-4o (Hurst et al., 2024). For the comparison with SVD, we use GPT-4o to generate five $16:9$ images for each prompt, which are resized to $1024\times 576$ as input. For the comparison with HunyuanVideo, we first use their released model to generate five videos at a resolution of $1280\times 720$ for each prompt, then extract the first frame of each video as the input image for our model to generate the corresponding video. No target pose is provided for any model. For each image, we generate one video per model using the same random seed ($42$), resulting in a total of $5,000$ video pairs. Each video pair is evaluated by three randomly selected users on Amazon MTurk, leading to a total of $15,000$ comparisons. Users are shown two videos side by side, generated by different models, with the order randomized. They are instructed to assess the videos based on Motion Consistency, Amount of Motion, and Motion Realism. The results are reported in Fig. 5. Our model significantly enhances the motion generation capabilities of SVD, producing videos with superior motion quantity, consistency, and realism. Furthermore, it even surpasses HunyuanVideo, a state-of-the-art video generation model with 13B parameters, in terms of motion quality. These results highlight the effectiveness of our model in generating complex motions and interactions.
@@
Qualitative Comparisons. We provide samples of generated videos in Fig. 4. Our ReVision produces realistic movements that closely follow user instructions. It also generates high-quality videos that involves complex motions and interactions, such as running with dogs, picking up a ball, and hitting a tennis ball. More visualizations are available in the supplementary videos.
Table 3: Inference speed. Average time to generate a 32-frame video. Our ReVision-SVD matches SVD in speed (8.4x faster than HunyuanVideo) while surpassing HunyuanVideo in generating complex motions and interactions.
@@
Time (s) 36 49 (8 + 5 + 36) 411
Inference Speed. We compare the inference speed of our model against two baselines in Tab. 3. Despite the three-step pipeline, the coarse video generation (S1) takes only 8 seconds after our optimization, and the additional 3D detection and refinement modules (S2) add 5 seconds to the inference time on a single A100. Together, these two stages are significantly faster than the original SVD, which requires 36 seconds. More importantly, with just 1.5B parameters and a runtime of 49 seconds, our model generates high-quality videos with complex motions – comparable to or even surpassing state-of-the-art models like HunyuanVideo (see Fig. 4 and  5), which uses over 13B parameters and requires an average of 411 seconds.
Figure 6: Handling occlusion. As illustrated by the two apples falling into the basket, ReVision handles occlusions by lifting and optimizing motion in 3D space, which allows explicit reasoning about object spatial relationships, effectively resolving occlusions that are ambiguous in 2D.
Handling Occlusion. Occlusion becomes a significant challenge when generating videos with multiple objects and large motions. However, by lifting everything into 3D, occlusion is naturally resolved: Since we estimate depth, all objects are fully represented with their spatial positions, allowing us to reason about their relative locations in 3D. And after optimizing the motion in 3D, we project it back to 2D using the depth information, which restores accurate occlusion relationships in the camera coordinate. An example is shown in Fig. 6, where two apples are generated dropping into a basket. Our model effectively captures spatial relationships, producing realistic videos in which the apples fall into the basket with appropriate occlusion.
Table 4: User preference comparisons for occlusion and interaction handling.
ReVision-SVD (Ours) SVD
Preference 97.63% 2.37%
ReVision-SVD (Ours) HunyuanVideo
Preference 63.99% 36.01%
To further evaluate our model’s ability to handle occlusion, we conducted an additional user study using the 5,000 video pairs generated for the main experiment. Users on Amazon MTurk were asked to assess the quality of occlusion handling and object interactions for each video pair. If no object interaction was observed, users were instructed to select "no interaction/occlusion." Similarly, each video pair was evaluated independently by three randomly selected users to ensure reliability. The study yielded 15,000 evaluations, including 10,207 valid comparisons and 4,793 responses marked as "no interaction/occlusion". The results of these 10,207 valid comparisons are summarized in Tab. 4, highlighting that our model consistently outperforms both SVD and HunyuanVideo across a diverse range of occlusion scenarios.
Figure 7: Long video generation. Our PMP extends a 32-frame 3D motion to 128 frames through interpolation (32 $\rightarrow$ 64), extrapolation (64 $\rightarrow$ 128), and refinement, enabling complex, large-scale motion generation over long video sequences. See supplementary videos for details.
Long Video Generation. Another advantage of our model is its ability to generate complex and large-scale motions over long video sequences (Fig. 7). PMP optimizes motion in a 3D parameterized space, enabling smooth and realistic interpolation and extrapolation to arbitrary lengths. The resulting long 3D motion sequences are then used to generate multiple overlapping video clips, which are stitched together to form extended videos with consistent motion. More specifically, we resample in motion-parameter space (not pixel space) through interpolation and extrapolation. Then, for long-video synthesis we generate overlapping 32-frame clips with a sliding window (stride 24, overlap 8) and stitch overlaps with a ramped blend in the overlapping frames. Motion conditions are aligned in the overlap and all windows share the same appearance/identity conditioning. The 3D representation plays a key role in maintaining smooth temporal continuity. While long video generation is not our main focus, using advanced techniques like temporal compression (Bar-Tal et al., 2024), beyond simple overlap, could further improve visual coherence. We leave their integration with our 3D-aware framework as future work.